If Henry did 8 problems for homework.If 2 were wrong, what percent of problems were wrong?
Answer by stanbon(75887)

If Henry did 8 problems for homework.If 2 were wrong, what percent of problems were wrong?
------
Equation:
Solve:
x/100 = 2/8
---
x/100 = 1/4
x = 100/4
x = 25
____
Answer: 25%
=======================
Cheers,
Stan H.